A slack rope walker balances on a rope tied between two trees. For a short, he is not moving. His mass is 72 KG. The rope makes an angle of 22 degrees with the horizontal on either side . What is the magnitude of the Tension in the rope ?


1
Expert's answer
2021-02-01T08:07:25-0500

"\text{external forces acting on the system:}"

"F_{net}=mg\text{ the force of gravity}"

"T_1,T_2 \text{ directed in different directions along the rope}"

"\text{to the suspension points}"

"\text{due to the symmetry of the system, the values of }"

"T_1\text{ and }T_2 \text{are equal }T_1=T_2"

"\text{the system is in equilibrium therefore the projections of the forces on the axis are 0}"

"F_{net y}=T_{1y}+T_{2y}"

"F_{net y}=F_{net}"

"T_{1y}=T*\sin\alpha"

"T_{2y}=T*\sin\alpha"

"T=\frac{Fnet}{2*\sin{\alpha}}=\frac{mg}{2*\sin{\alpha}}=\frac{72*9.8}{2*\sin22^0}\approx 941.8"

Answer:941.8 N